Operation Summary for April 13, 2011 - April 15, 2011 |
|
Wednesday April 13 The day shift began with the Tevatron (TEV) going into a wet squeeze, with Operations monitoring the Recycler stash (312.0E10) and the Antiproton Source (Pbar) stack (19.6E10), with the Main Injector (MI) and Booster waiting for Linac beam, with Linac off due to problems with a Linac RF station (LRF1), with the NuMI and MiniBooNE waiting on beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, Operations reported that the TEV Proton Abort Kicker (PAK3) had prefired a number of times throughout the day in spite of lowered reservoir voltage. Kicker personnel decided that they had to open up the kicker supply tank and change out the tube. Pbar personnel accessed their ring and pre-vault area from 10:26 AM to 1 PM. Operators prepared the TEV A-E enclosures and the D0 collision hall for access. FESS personnel changed the MI-52 pond pump to its high-speed summer mode. Linac and Booster personnel conducted Neutron Therapy Facility (NTF) beam studies at various times during this shift. NuMI personnel reported that their newly installed target had developed a LCW leak. In order to keep this leak under control, Operations lowered the NuMI beam intensity to 1.5E13. Operations reported that they had stashed antiprotons until they reached the stash limit of 330E10; further antiprotons were held in the Pbar Accumulator. The Meson FTFB T-978 experiment did not take any beam due to problems with equipment in their experimental hall. During the evening shift, TEV PAK3 repairs continued. Booster personnel tuned to correct injection beam positions and to increase beam intensity to MiniBooNE. Operations reported that antiprotons were stacked and held in the Accumulator. Thursday April 14 The midnight shift began with the TEV still off due to kicker problems, with Operations monitoring the stash (325.8E10), with the Pbar stacking (172.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, there was no TEV beam due to continuing PAK3 work. Operations established beam to T-978 at 4:44 AM. Operations conducted Booster and MI tuning. Pbar stacked antiprotons and held them in the Accumulator. During the day shift, PAK3 work continued. MTA took beam from 10:08 AM to 12:33 PM. Operations established a TEV QBS ramp at 10:37 AM. Operations completed the TEV main bus checkout at 10:58 AM. Operations reported that a TEV RF station (TRF2) power amplifier (PA) failed; RF technicians had the PA replaced by 2:10 PM. C0 cryo system work was also completed at 2:10 PM. Operations put the TEV through a dry squeeze. TEV personnel conducted proton studies. Pbar stacked antiprotons and held them in the Accumulator. During the evening shift, TEV personnel conducted damper studies until 6:07 PM. Operations reported that the kicker techs had completed the TEV abort kicker repairs at 9:53 PM; Operations put the kickers in conditioning mode for the rest of the shift. Operations prepared D0 for an access. Operations disabled beam to T-978 at 8 PM. MTA took beam for a short time. Bad Booster beam positions caused some downtime. Pbar stacked antiprotons and held them in the Accumulator. Friday April 15 The midnight shift began with the TEV abort kickers in conditioning mode, with Operations monitoring the stash (318.2E10), with Pbar stacking (283.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations prepared D0 for an access early in the shift. Operations reported at 3 AM that the TEV abort kicker conditioning had been completed. Operations put the TEV through a series of one-hour dry squeezes. Operations reported that a bug in the antiproton flusher program caused the loss of about 50mA of antiprotons from the stack. Pbar stacked antiprotons and held them in the Accumulator. The Plans for Friday and the Weekend The plan for today is to put the TEV into shot setup as soon as possible after TEV personnel complete their kicker studies. The plan for the weekend is to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 04/15/2011 email Fermilab |
| Security, Privacy, Legal | |